Superheated vapor tables show that at 5500 psia, 1200°F the density of water vapor is 6.6 lb/ft³. Determine the percentage error involved if the density is calculated by using (a) the modified reduced van der Waals equation, (b) the ideal gas theory.

1
Expert's answer
2021-12-14T09:28:47-0500

"\rho=\frac{pM}{RT}=5.4~\frac{lb}{ft^3},"

"\eta=(1-\frac{\rho}{\rho_0})\cdot 100\%=18\%."